HIGH RESOLUTION SPECTROSCOPY OF Ar-CH_{4} IN THE 7 μ REGION

\underline{I. PAK}, M. HEPP, D. A. ROTH, G. WINNEWISSER, I. Physikalisches Institut, Universität zu Köln, 50937 Köln, Germany; D. SCOUTERIS, B. J. HOWARD, Physical Chemistry Laboratory, Oxford University, South Parks Road, Oxford OX1 3QZ, United Kingdom; K. M. T. YAMADA, National Institute for Advanced Interdisciplinary Research (NAIR), Higashi 1-1-4, Tsukuba, Ibaraki 305, Japan.

The absorption spectrum of the Ar-CH_{4} van der Waals complexes was measured in the 7 μ region. The complexes were produced in a pulsed slit supersonic jet and detected by a tunable diode laser spectrometer with 40 MHz spectral resolution. The strongest absorption features of the Ar-CH_{4} were observed in the spectral region 1310.5–1311.1 cm^{−1} near the R(0) line of the ν_{4} fundamental band of the methane monomer. The observed transitions are related to the j = 1 ← 0 transition of the CH_{4} in the complex. Analysis of the data is in progress, and the results of the analysis will be presented at the conference.